OpenStudy (anonymous):

x^2+y^2+x+y-1/2=0 find the center(h,k) and radius r of each circle(B) Graph each circle(c)find the intercepts...

OpenStudy (anonymous):

(x + 1/2)^2 - 1/4 + (y + 1/2) - 1/4 = 1/2 (x+ 1/2)^2 + (y+1/2)^2 = 1 center is (-1/2, -1/2) and radius = 1

OpenStudy (anonymous):

wow!!! I was not even close ..... thank you so much

OpenStudy (anonymous):

theres only one circle

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

and that's the -1/2, -1/2

OpenStudy (anonymous):

yes thats the coordinates of the center do you need the x and y intercepts as well?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I am terrible at math. taking a college pre calculus and feeling really dumb right notw...lol

OpenStudy (anonymous):

(x+ 1/2)^2 + (y+1/2)^2 = 1 when y = 0 (x+ 1/2)^2 + (1/2)^2 = 1 x^2 + x + 1/4 + 1/4 = 1 x^2 + x - 0.5 = 0 x = 0.366 and - 1.366 = x- intercepts

OpenStudy (anonymous):

y intercepts have the same values

OpenStudy (anonymous):

best way to improve is to practice problems as often as u can - good luck

OpenStudy (anonymous):

I will ... thank you so much